OpenStudy (anonymous):

Translate to a system of equations do not solve Tyler is twice as old as his son. Ten years ago, tyler was three times as old as his son. How old are they?

OpenStudy (welshfella):

let tylers age be x and his sons age be y now as tyler is 2* his sons age how do we write this as an equation?

OpenStudy (anonymous):

t = 2s --- It says that Tyler is twice as old as his son. t - 10 = 3(s -10) --- This is because if you take t, Tyler's current age, and subtract 10 from it, you get Tyler's age 10 years ago. However, you must also take those 10 years away from his son, because they would have both been 10 years younger. After that, you just need to multiple his son's age by 3 and set the two expressions equal to each other. Hope I helped

OpenStudy (welshfella):

s = something times y

OpenStudy (welshfella):

* x = something * y

OpenStudy (welshfella):

what is that something?

OpenStudy (mertsj):

t = tyler's age s = son's age t=2s TEN YEARS AGO: tyler's age: t-10 son's age: s-10 t-10=3(s-10)